OpenStudy (anonymous):

lim cos2x - cos 3x/cos4x - 1 x--> 0

OpenStudy (ipwnbunnies):

Just to be sure, is it this? \[\lim_{x \rightarrow 0}\frac{\cos 2x - \cos 3x}{\cos 4x - 1}\]

OpenStudy (anonymous):

that's the same thing I got also

OpenStudy (anonymous):

yup, I need to evalute the limit

OpenStudy (ipwnbunnies):

The first thing you should try is direct substitution: Plugging in 0 for x.

OpenStudy (anonymous):

eh.. it direct submission wouldn't make everything zero right?

OpenStudy (ipwnbunnies):

What do you mean? :3 Plug in 0 for x and evaluate. What do you get?

OpenStudy (anonymous):

@iPwnBunnies Plug in 0 for x, we'll get \(\frac{0}{0}\).

OpenStudy (anonymous):

if you have \(\frac{0}{0}\) or \(\frac{\infty}{\infty}\), where are indeterminate, you can use l'Hopital's rule. \[\lim_{x\rightarrow0}\frac{\cos(2x)-\cos(3x)}{\cos(4x)-1}\]\[=\lim_{x\rightarrow0}\frac{[\cos(2x)-\cos(3x)]'}{[\cos(4x)-1]'}\]\[=...\]That ' is differentiation. In this question, you need to apply this rule twice.